A beam of light is incident from air on the surface of a liquid. If the angle of incidence is 29.5
Answer #1

refractive index = sin29.5 / sin17.1 = 1.675

so critical angle = sin-1(1/1.675) = 36.66 degrees
Answer #2

sin29.5 = u*sin17.1

u = 1.67

critical angle:

1.67sin(C) = 1

c = 36.78 degree